OpenStudy (dragon123mama):

how much money invested at 5% compounded contentiously for 3 years will result in $820?

OpenStudy (dragon123mama):

continuously**

pooja195 (pooja195):

Any ideas on where to start?

OpenStudy (raffle_snaffle):

Use F = (1+P)^n where P is the present cost, F is the final cost, and n is the # of years.

OpenStudy (dragon123mama):

okay then what?

pooja195 (pooja195):

Plug in the info into the formula

OpenStudy (dragon123mama):

??

pooja195 (pooja195):

\(\color{#0cbb34}{\text{Originally Posted by}}\) @raffle_snaffle Use F = (1+P)^n where P is the present cost, F is the final cost, and n is the # of years. \(\color{#0cbb34}{\text{End of Quote}}\) Thise

OpenStudy (dragon123mama):

yeah i saw that. but where did it come from?

pooja195 (pooja195):

That is the formula we would use to find out how much money is being invested

OpenStudy (dragon123mama):

oh. i didn't know that.

OpenStudy (dragon123mama):



OpenStudy (dragon123mama):

would that be it?

pooja195 (pooja195):

\[\huge~\rm~\bf~A=P(1+\frac{ r }{ 100 })^n\] We know: A=820 n=3 r=5 Now you just have to find P So plug in the given info into that equation

OpenStudy (dragon123mama):

what. you told me to use that first equation. i am confused. what is this second equation have to do with anything.

pooja195 (pooja195):

We use the compound interest formula to find investment the other user made a mistake as did I, I apologize.

OpenStudy (dragon123mama):

okay.... so what is this second equation?

OpenStudy (dragon123mama):

i don't know how to use that one.

pooja195 (pooja195):

Compound interest formula plug in the info I gave you and you basically solve for P .

OpenStudy (dragon123mama):

i don't know how to solve that complex of an equation.

OpenStudy (dragon123mama):

my calculator broke otherwise i would just use that to solve it.

pooja195 (pooja195):

\(\color{#0cbb34}{\text{Originally Posted by}}\) @pooja195 \[\huge~\rm~\bf~A=P(1+\frac{ r }{ 100 })^n\] We know: A=820 n=3 r=5 Now you just have to find P So plug in the given info into that equation \(\color{#0cbb34}{\text{End of Quote}}\) \[\huge~\rm~\bf~820=P(1+\frac{ 5 }{ 100 })^3\] Solve for p.. give it a try

OpenStudy (dragon123mama):

i can't do that without a calculator. it is way to complicated to do without one.

OpenStudy (dragon123mama):

i don't know how to solve this complex on paper.

OpenStudy (dragon123mama):

why are there so many people on my post that aren't even doing anything. if you can't help on this question plz go help on my other one.

pooja195 (pooja195):

If you aren't gonna try and just troll there isn't really a point in helping you.

OpenStudy (mathstudent55):

You are dealing with continuous compounding. You need the formula that deals with that. It uses the constant e.

OpenStudy (dragon123mama):

im seriously not. my calculator broke. and i didn't learn how to solve these types of things on paper. iv always just used a calculator. i am not trying to troll i sware.

OpenStudy (wolf1728):

dragon 1) you need a CONTINUOUS compounding formula and 2) I am looking for the formula to post here.

OpenStudy (mathstudent55):

\(\Large A = Pe^{rt} \)

OpenStudy (dragon123mama):

what. more formulas. now i am more confused oh my.

OpenStudy (dragon123mama):

how do i solve that formula now then if i need that one?

pooja195 (pooja195):

You would get the same answer with either formula.

OpenStudy (mathstudent55):

You have the end amount, A. You are looking for P. e is a famous math constant. r = 5% = 0.05 t = 3

OpenStudy (wolf1728):

Here it is:

OpenStudy (mathstudent55):

Continuous compounding yields more interest than periodic compounding.

OpenStudy (dragon123mama):

okay. so then what?

OpenStudy (dragon123mama):

iv never even seen this formula before.

OpenStudy (mathstudent55):

\(\Large A = Pe^{rt} \) \(\Large 820 = Pe^{0.05 \times 3} \) \(\Large P = \dfrac{820}{e^{0.15}} \)

OpenStudy (wolf1728):

5% 3 years $820 Principal = Total / e^[rate * years] Principal = 820 / e^.15 (which is what mathstudent just posted)

OpenStudy (dragon123mama):

okay. so that is the answer?

OpenStudy (mathstudent55):

Use a calculator. A scientific calculator can raise e to an exponent.

OpenStudy (dragon123mama):

my calculator broke. well the batteries died i don't have money for more.

pooja195 (pooja195):

I think you've trolled enough. Good bye.

OpenStudy (dragon123mama):

scientific calculator is so expensive i have never owned one

OpenStudy (sachintha):

Use google. It will solve it for you. :3

OpenStudy (mathstudent55):

Here is a free one. https://www.google.com/#q=scientific+calculator

OpenStudy (dragon123mama):

i am not trolling. these people are helping me lots. it is starting to make sence. you havent helpes at all pooja

OpenStudy (dragon123mama):

thanks guys. i didn't know there was one online. give me a sec then.

OpenStudy (wolf1728):

No you have to do the computing. but your calculator is broken? dragon here is an online calculator : http://www.1728.org/calc.htm it WILL raise numbers to a power

OpenStudy (dragon123mama):

705.780540669 this was the answer it showed. this correct?

OpenStudy (wolf1728):

That's not the answer I got.

OpenStudy (mathstudent55):

Correct.

OpenStudy (dragon123mama):

it is what the calculator online showed.

OpenStudy (dragon123mama):

although i don't think you can invest money with so many numbers after that decimal point.

OpenStudy (wolf1728):

Sorry, I recalculated. That answer is correct.

OpenStudy (mathstudent55):

If you do annual compounding, then the answer is $708.35 It makes sense that if you do annual compounding you have to start with a little more principal because continuous compounding yields more interest that periodic compounding.

OpenStudy (dragon123mama):

oh okay thanks. can you guys help with my question in mathematics please?